# Adaptive Prompt: Unlocking the Power of Visual Prompt Tuning

### v1

@@ Appendix B Additional Theoretical Results for Visual Adaptive Prompt Tuning
Here, we denote $e_{u}$ is the vector in $\mathbb{R}^{d}$ such that its $u$-th element is 1 while its other elements are 0 for any $1\leq u\leq d$, while $e_{uv}$ stands for the matrix in $\mathbb{R}^{d\times d}$ with its $uv$-th entry is 1 while other entries are zero.
Decomposition of $B_{n}({\bm{X}})$.
We can rewrite $B_{n}({\bm{X}})$ as follows:
$B_{n}({\bm{X}})$ $=\sum_{j:|\mathcal{V}_{j}|=1}\sum_{i\in\mathcal{V}_{j}}\exp(b_{n,i})\Big{[}E({\bm{X}};W_{n,2}W_{n,1i})-E({\bm{X}};W_{*,2}W_{*,1j})\Big{]}f_{G_{n}}({\bm{X}})$
$+\sum_{j:|\mathcal{V}_{j}|>1}\sum_{i\in\mathcal{V}_{j}}\exp(b_{n,i})\Big{[}E({\bm{X}};W_{n,2}W_{n,1i})-E({\bm{X}};W_{*,2}W_{*,1j})\Big{]}f_{G_{n}}({\bm{X}})$
$:=B_{n,1}({\bm{X}})+B_{n,2}({\bm{X}}).$
By applying the first-order and second-order Taylor expansions, we get
$B_{n,1}({\bm{X}})$ $=\sum_{j:|\mathcal{V}_{j}|=1}\sum_{|\alpha|=1}M_{n,j,\alpha}\dfrac{\partial^{|\alpha|}E}{\partial{(W_{2}W_{1})^{\alpha}}}({\bm{X}};W_{*,2}W_{*,1j})f_{G_{n}}({\bm{X}})+R_{n,3}({\bm{X}}),$
$B_{n,2}({\bm{X}})$ $=\sum_{j:|\mathcal{V}_{j}|=1}\sum_{1\leq|\alpha|\leq 2}M_{n,j,\alpha}\dfrac{\partial^{|\alpha|}E}{\partial{(W_{2}W_{1})^{\alpha}}}({\bm{X}};W_{*,2}W_{*,1j})f_{G_{n}}({\bm{X}})+R_{n,4}({\bm{X}})$
where $R_{n,3}({\bm{X}}),R_{n,4}({\bm{X}})$ is a Taylor remainder such that $R_{n,3}({\bm{X}})/\mathcal{D}_{2n}\to 0$, $R_{n,4}({\bm{X}})/\mathcal{D}_{2n}\to 0$ when $n\to\infty$. Therefore, we can express the functions $B_{n,1}({\bm{X}})$ and $B_{n,2}({\bm{X}})$ as follows:
$B_{n,1}({\bm{X}})=\sum_{j:|\mathcal{V}_{j}|=1}\exp((W_{*,2}W_{*,1j}{\bm{X}})^{\top}{\bm{X}})\Big{[}\sum_{u_{1},v_{1}=1}^{d}M_{n,j,e_{u_{1}v_{1}}}{\bm{X}}^{(u_{1})}{\bm{X}}^{(v_{1})}\Big{]}f_{G_{n}}({\bm{X}})+{R}_{n,3}({\bm{X}}),$
$B_{n,2}({\bm{X}})=\sum_{j:|\mathcal{V}_{j}|>1}\exp((W_{*,2}W_{*,1j}{\bm{X}})^{\top}{\bm{X}})\Big{[}\sum_{u_{1},v_{1}=1}^{d}M_{n,j,e_{u_{1}v_{1}}}{\bm{X}}^{(u_{1})}{\bm{X}}^{(v_{1})}$
$\hskip 113.81102pt+\sum_{u_{1},v_{1}=1}^{d}\sum_{u_{2},v_{2}=1}^{d}M_{n,j,e_{u_{1}v_{1}}}{\bm{X}}^{(u_{1})}{\bm{X}}^{(v_{1})}{\bm{X}}^{(u_{2})}{\bm{X}}^{(v_{2})}\Big{]}f_{G_{n}}({\bm{X}})+R_{n,4}({\bm{X}}),$
Plugging the above expressions into equation (32), we can represent $Q_{n}({\bm{X}})$ as follows:
$Q_{n}({\bm{X}})$ $=\sum_{j:|\mathcal{V}_{j}|=1}\exp((W_{*,2}W_{*,1j}{\bm{X}})^{\top}{\bm{X}})\Big{[}\sum_{u_{1},v_{1}=1}^{d}M_{n,j,e_{u_{1}v_{1}}}{\bm{X}}^{(u_{1})}{\bm{X}}^{(v_{1})}W_{*,2}W_{*,1j}{\bm{X}}+\sum_{u_{1},v_{1}=1}^{d}M_{n,j,e_{u_{1}v_{1}}}{\bm{X}}^{(v_{1})}e_{u_{1}}\Big{]}$
$+\sum_{j:|\mathcal{V}_{j}|>1}\exp((W_{*,2}W_{*,1j}{\bm{X}})^{\top}{\bm{X}})\Big{[}\sum_{u_{1},v_{1}=1}^{d}M_{n,j,e_{u_{1}v_{1}}}{\bm{X}}^{(u_{1})}{\bm{X}}^{(v_{1})}W_{*,2}W_{*,1j}{\bm{X}}+\sum_{u_{1},v_{1}=1}^{d}M_{n,j,e_{u_{1}v_{1}}}{\bm{X}}^{(v_{1})}e_{u_{1}}$
$\hskip 56.9055pt+\sum_{u_{1},v_{1}=1}^{d}\sum_{u_{2},v_{2}=1}^{d}M_{n,j,e_{u_{1}v_{1}}+e_{u_{2}v_{2}}}{\bm{X}}^{(u_{1})}{\bm{X}}^{(v_{1})}{\bm{X}}^{(u_{2})}{\bm{X}}^{(v_{2})}W_{*,2}W_{*,1j}{\bm{X}}$
$\hskip 56.9055pt+\sum_{u_{1},v_{1}=1}^{d}\sum_{u_{2},v_{2}=1}^{d}M_{n,j,e_{u_{1}v_{1}},e_{u_{2}v_{2}}}{\bm{X}}^{(u_{1})}{\bm{X}}^{(v_{1})}{\bm{X}}^{(v_{2})}e_{u_{2}}\Big{]}$
$-\sum_{j:|\mathcal{V}_{j}|=1}\exp((W_{*,2}W_{*,1j}{\bm{X}})^{\top}{\bm{X}})\Big{[}\sum_{u_{1},v_{1}=1}^{d}M_{n,j,e_{u_{1}v_{1}}}{\bm{X}}^{(u_{1})}{\bm{X}}^{(v_{1})}\Big{]}f_{G_{n}}({\bm{X}})$
$-\sum_{j:|\mathcal{V}_{j}|>1}\exp((W_{*,2j}W_{*,1j}{\bm{X}})^{\top}{\bm{X}})\Big{[}\sum_{u_{1},v_{1}=1}^{d}M_{n,j,e_{u_{1}v_{1}}}{\bm{X}}^{(u_{1})}{\bm{X}}^{(v_{1})}$
$\hskip 113.81102pt+\sum_{u_{1},v_{1}=1}^{d}\sum_{u_{2},v_{2}=1}^{d}M_{n,j,e_{u_{1}v_{1}}}{\bm{X}}^{(u_{1})}{\bm{X}}^{(v_{1})}{\bm{X}}^{(u_{2})}{\bm{X}}^{(v_{2})}\Big{]}f_{G_{n}}({\bm{X}})$
$-\sum_{j=1}^{L}N_{n,j}\exp((W_{*,2}W_{*,1j}{\bm{X}})^{\top}{\bm{X}})f_{G_{n}}({\bm{X}})$
$+\sum_{j=1}^{L}N_{n,j}\exp((W_{*,2}W_{*,1j}{\bm{X}})^{\top}{\bm{X}})W_{*,2}W_{*,1j}{\bm{X}}$
$+{R}_{n,1}({\bm{X}})+{R}_{n,2}({\bm{X}})-R_{n,3}({\bm{X}})-R_{n,4}({\bm{X}})$ (33)
where $N_{n,j}:=\sum_{i\in\mathcal{V}_{j}}\exp(b_{n,i})-\exp(b_{*,j})$ for any $j\in[L]$.
@@
From equation (33), we can represent $Q_{n}({\bm{X}})/\mathcal{D}_{2n}$ as a linear combination of the following independent functions:
@@
Assume that all the coefficients of these linear independent functions in the formulation of $Q_{n}({\bm{X}})/\mathcal{D}_{2n}$ go to 0 as $n\to\infty$. It follows that $M_{n,j,\alpha}/\mathcal{D}_{2n}$, $M_{n,j,\alpha,\beta}/\mathcal{D}_{2n}$, and $N_{n,j}/\mathcal{D}_{2n}$ approach 0 as $n\to\infty$, for all $\alpha,\beta\in\mathbb{N}^{d\times d}$ such that $1\leq|\alpha|+|\beta|\leq 2$.
Then, as $N_{n,j}/\mathcal{D}_{2n}\to 0$, it indicates that
$\frac{|\sum_{i\in\mathcal{V}_{j}}\exp(b_{n,i})-\exp(b_{*,j})|}{\mathcal{D}_{2n}}=\frac{|N_{n,j}|}{\mathcal{D}_{2n}}\to 0,$
for any $1\leq j\leq L$. By summing these limits up when varying the index $j$ from 1 to $L$, we obtain that
@@
Now, we consider indices $j\in[L]$ such that its corresponding Voronoi cell has only one element, i.e. $|\mathcal{V}_{j}|=1$. As $M_{n,j,e_{uv}}/\mathcal{D}_{2n}\to 0$, we have
$\frac{\sum_{i\in\mathcal{V}_{j}}\exp(b_{n,i})\|\Delta W_{n,2}W_{n,1ij}\|_{1}}{\mathcal{D}_{2n}}=\frac{\sum_{u,v=1}^{d}|M_{n,j,e_{uv}}|}{\mathcal{D}_{2n}}\to 0.$
@@
Moving to indices $j\in[L]$ such that $|\mathcal{V}_{j}|>1$, as $M_{n,j,2e_{uv}}/\mathcal{D}_{2n}\to 0$, we obtain that
$\frac{\sum_{i\in\mathcal{V}_{j}}\exp(b_{n,i})\|\Delta W_{n,2}W_{n,1ij}\|^{2}}{\mathcal{D}_{2n}}=\frac{\sum_{u,v=1}^{d}M_{n,j,2e_{uv}}}{\mathcal{D}_{2n}}\to 0.$ (36)
Collecting all the results in equations (34), (35), and (36), we obtain that
$1=\frac{\mathcal{D}_{2n}}{\mathcal{D}_{2n}}\to 0$
as $n\to\infty$, which is a contradiction.
As a consequence, not all the coefficients of the linear independent functions in the formulations of $Q_{n}({\bm{X}})/\mathcal{D}_{2n}$ go to 0 as $n\to\infty$.
@@
Let us denote $m_{n}$ as the maximum of the absolute values of $M_{n,j,\alpha}/\mathcal{D}_{2n}$, $M_{n,j,\alpha,\beta}/\mathcal{D}_{2n}$, and $N_{n,j}/\mathcal{D}_{2n}$ for all $\alpha,\beta\in\mathbb{N}^{d\times d}$ such that $1\leq|\alpha|+|\beta|\leq 2$. From the result of Step 2, it follows that $1/m_{n}\not\to\infty$ as $n\to\infty$. Since $\|f_{G_{n}}-f_{G_{*}}\|_{L_{2}(\mu)}/\mathcal{D}_{2n}\to 0$ as $n\to\infty$, we obtain $\|f_{G_{n}}-f_{G_{*}}\|_{L_{2}(\mu)}/(m_{n}\mathcal{D}_{2n})\to 0.$ By applying Fatou’s lemma, we get that
$0=\lim_{n\to\infty}\dfrac{\|f_{G_{n}}-f_{G_{*}}\|_{L_{2}(\mu)}}{m_{n}\mathcal{D}_{2n}}\geq\int\liminf_{n\to\infty}\dfrac{\left|f_{G_{n}}({\bm{X}})-f_{G_{*}}({\bm{X}})\right|}{m_{n}\mathcal{D}_{2n}}d\mu({\bm{X}})\geq 0$
Therefore, $\liminf_{n\to\infty}\dfrac{\left|f_{G_{n}}({\bm{X}})-f_{G_{*}}({\bm{X}})\right|}{m_{n}\mathcal{D}_{2n}}=0$ for almost surely ${\bm{X}}$. As $n\to\infty$, we denote
$\dfrac{M_{n,j,\alpha}}{m_{n}\mathcal{D}_{2n}}\to\lambda_{j,\alpha},\quad\dfrac{M_{n,j,\alpha,\beta}}{m_{n}\mathcal{D}_{2n}}\to\xi_{j,\alpha,\beta},\quad\dfrac{N_{n,j}}{m_{n}\mathcal{D}_{2n}}\to\tau_{j},$
for any $1\leq j\leq L$ and $\alpha,\beta\in\mathbb{N}^{d\times d}$ such that $1\leq|\alpha|+|\beta|\leq 2$. Here, from the definition of $m_{n}$, at least one coefficient among $\{\lambda_{j,\alpha},\xi_{j,\alpha,\beta},\tau_{j}:j\in[L],\alpha,\beta\in\mathbb{N}^{d\times d}:1\leq|\alpha|+|\beta|\leq 2\}$ is different from 0. Then, the equation $\liminf_{n\to\infty}\dfrac{\left|f_{G_{n}}({\bm{X}})-f_{G_{*}}({\bm{X}})\right|}{m_{n}\mathcal{D}_{2n}}=0$, or equivalently, $\liminf_{n\to\infty}\dfrac{\left|Q_{n}({\bm{X}})\right|}{m_{n}\mathcal{D}_{2n}}=0$ leads to
@@
$\lim_{\varepsilon\to 0}\inf_{G\in\mathcal{G}_{L^{\prime}}(\Theta):\mathcal{D}_{2}(G,G_{*})\leq\varepsilon}\|f_{G}-f_{G_{*}}\|_{L_{2}(\mu)}/\mathcal{D}_{2}(G,G_{*})>0.$
Global part:
@@
From local part, there exists a positive constant $\varepsilon^{\prime}$ such that
$\inf_{G\in\mathcal{G}_{L^{\prime}}(\Theta):\mathcal{D}_{2}(G,G_{*})\leq\varepsilon^{\prime}}\|f_{G}-f_{G_{*}}\|_{L_{2}(\mu)}/\mathcal{D}_{2}(G,G_{*})>0.$
Therefore, it is sufficient to prove that
$\inf_{G\in\mathcal{G}_{L^{\prime}}(\Theta):\mathcal{D}_{2}(G,G_{*})>\varepsilon^{\prime}}\|f_{G}-f_{G_{*}}\|_{L_{2}(\mu)}/\mathcal{D}_{2}(G,G_{*})>0.$
Assume by contrary, then we can find a sequence of mixing measures $G^{\prime}_{n}:=\sum_{j^{\prime}=1}^{L^{\prime}}\exp(b_{n,j^{\prime}})\delta_{W_{n,2}W_{n,1j^{\prime}}}$ in $\mathcal{G}_{L^{\prime}}(\Theta)$ such that as $n\to\infty$, we have

### v3

@@ Appendix B Additional Theoretical Results for Visual Adaptive Prompt Tuning
Here, we denote $e_{u}$ is the standard basis vector in $\mathbb{R}^{d}$ with 1 in the $u$-th position and 0 in all other positions for any $1\leq u\leq d$, while $e_{uv}$ stands for the matrix in $\mathbb{R}^{d\times d}$ with its $uv$-th entry is 1 while other entries are zero.
Decomposition of the function $\tilde{B}_{n}({\bm{X}})$.
Similar to our decomposition of the function $\tilde{A}_{n}({\bm{X}})$, we also separate the Voronoi cells with exactly one element and with more than one element in the decomposition of the function $\tilde{B}_{n}({\bm{X}})$. Therefore, we obtain that
$\tilde{B}_{n}({\bm{X}})$ $=\sum_{j:|\mathcal{V}_{j}|=1}\sum_{i\in\mathcal{V}_{j}}\exp(b_{n,i})\Big{[}\tilde{E}({\bm{X}};W_{n,2}W_{n,1i})-\tilde{E}({\bm{X}};W_{*,2}W_{*,1j})\Big{]}f_{G_{n}}({\bm{X}})$
$+\sum_{j:|\mathcal{V}_{j}|>1}\sum_{i\in\mathcal{V}_{j}}\exp(b_{n,i})\Big{[}\tilde{E}({\bm{X}};W_{n,2}W_{n,1i})-\tilde{E}({\bm{X}};W_{*,2}W_{*,1j})\Big{]}f_{G_{n}}({\bm{X}})$
$:=\tilde{B}_{n,1}({\bm{X}})+\tilde{B}_{n,2}({\bm{X}}).$
For Voronoi cells with exactly one element, we will utilize Taylor expansion up to the first order while for Voronoi cells with more than one element, the second-order Taylor expansion will be employed. This strategy leads to the following representations:
$\tilde{B}_{n,1}({\bm{X}})$ $=\sum_{j:|\mathcal{V}_{j}|=1}\sum_{|\alpha|=1}\tilde{M}_{n,j,\alpha}\dfrac{\partial^{|\alpha|}\tilde{E}}{\partial{(W_{2}W_{1})^{\alpha}}}({\bm{X}};W_{*,2}W_{*,1j})f_{G_{n}}({\bm{X}})+\tilde{R}_{n,3}({\bm{X}}),$
$\tilde{B}_{n,2}({\bm{X}})$ $=\sum_{j:|\mathcal{V}_{j}|=1}\sum_{1\leq|\alpha|\leq 2}\tilde{M}_{n,j,\alpha}\dfrac{\partial^{|\alpha|}\tilde{E}}{\partial{(W_{2}W_{1})^{\alpha}}}({\bm{X}};W_{*,2}W_{*,1j})f_{G_{n}}({\bm{X}})+\tilde{R}_{n,4}({\bm{X}}).$
In these expressions, the functions $\tilde{R}_{n,3}({\bm{X}})$ and $\tilde{R}_{n,4}({\bm{X}})$ correspond to the Taylor remainders. Due to the uniform smoothness of the functions $\tilde{E}$, we obtain that $\tilde{R}_{n,3}({\bm{X}})/\mathcal{D}_{2n}\to 0$ and $\tilde{R}_{n,4}({\bm{X}})/\mathcal{D}_{2n}\to 0$ as $n$ approaches infinity. By computing the closed-form formulations for the partial derivatives of the function $\tilde{E}$, both the functions $\tilde{B}_{n,1}({\bm{X}})$ and $\tilde{B}_{n,2}({\bm{X}})$ can be rewritten as follows:
$\tilde{B}_{n,1}({\bm{X}})=\sum_{j:|\mathcal{V}_{j}|=1}\exp((W_{*,2}W_{*,1j}{\bm{X}})^{\top}{\bm{X}})\Big{[}\sum_{u_{1},v_{1}=1}^{d}\tilde{M}_{n,j,e_{u_{1}v_{1}}}{\bm{X}}^{(u_{1})}{\bm{X}}^{(v_{1})}\Big{]}f_{G_{n}}({\bm{X}})+\tilde{R}_{n,3}({\bm{X}}),$
$\tilde{B}_{n,2}({\bm{X}})=\sum_{j:|\mathcal{V}_{j}|>1}\exp((W_{*,2}W_{*,1j}{\bm{X}})^{\top}{\bm{X}})\Big{[}\sum_{u_{1},v_{1}=1}^{d}\tilde{M}_{n,j,e_{u_{1}v_{1}}}{\bm{X}}^{(u_{1})}{\bm{X}}^{(v_{1})}$
$\hskip 113.81102pt+\sum_{u_{1},v_{1}=1}^{d}\sum_{u_{2},v_{2}=1}^{d}\tilde{M}_{n,j,e_{u_{1}v_{1}}}{\bm{X}}^{(u_{1})}{\bm{X}}^{(v_{1})}{\bm{X}}^{(u_{2})}{\bm{X}}^{(v_{2})}\Big{]}f_{G_{n}}({\bm{X}})+\tilde{R}_{n,4}({\bm{X}}),$
Plugging all of these results together, the function $Q_{n}({\bm{X}})$ can be rewritten as follows:
$Q_{n}({\bm{X}})$ $=\sum_{j:|\mathcal{V}_{j}|=1}\exp((W_{*,2}W_{*,1j}{\bm{X}})^{\top}{\bm{X}})\Big{[}\sum_{u_{1},v_{1}=1}^{d}\tilde{M}_{n,j,e_{u_{1}v_{1}}}{\bm{X}}^{(u_{1})}{\bm{X}}^{(v_{1})}W_{*,2}W_{*,1j}{\bm{X}}+\sum_{u_{1},v_{1}=1}^{d}\tilde{M}_{n,j,e_{u_{1}v_{1}}}{\bm{X}}^{(v_{1})}e_{u_{1}}\Big{]}$
$+\sum_{j:|\mathcal{V}_{j}|>1}\exp((W_{*,2}W_{*,1j}{\bm{X}})^{\top}{\bm{X}})\Big{[}\sum_{u_{1},v_{1}=1}^{d}\tilde{M}_{n,j,e_{u_{1}v_{1}}}{\bm{X}}^{(u_{1})}{\bm{X}}^{(v_{1})}W_{*,2}W_{*,1j}{\bm{X}}+\sum_{u_{1},v_{1}=1}^{d}\tilde{M}_{n,j,e_{u_{1}v_{1}}}{\bm{X}}^{(v_{1})}e_{u_{1}}$
$\hskip 56.9055pt+\sum_{u_{1},v_{1}=1}^{d}\sum_{u_{2},v_{2}=1}^{d}\tilde{M}_{n,j,e_{u_{1}v_{1}}+e_{u_{2}v_{2}}}{\bm{X}}^{(u_{1})}{\bm{X}}^{(v_{1})}{\bm{X}}^{(u_{2})}{\bm{X}}^{(v_{2})}W_{*,2}W_{*,1j}{\bm{X}}$
$\hskip 56.9055pt+\sum_{u_{1},v_{1}=1}^{d}\sum_{u_{2},v_{2}=1}^{d}\tilde{M}_{n,j,e_{u_{1}v_{1}},e_{u_{2}v_{2}}}{\bm{X}}^{(u_{1})}{\bm{X}}^{(v_{1})}{\bm{X}}^{(v_{2})}e_{u_{2}}\Big{]}$
$-\sum_{j:|\mathcal{V}_{j}|=1}\exp((W_{*,2}W_{*,1j}{\bm{X}})^{\top}{\bm{X}})\Big{[}\sum_{u_{1},v_{1}=1}^{d}\tilde{M}_{n,j,e_{u_{1}v_{1}}}{\bm{X}}^{(u_{1})}{\bm{X}}^{(v_{1})}\Big{]}f_{G_{n}}({\bm{X}})$
$-\sum_{j:|\mathcal{V}_{j}|>1}\exp((W_{*,2j}W_{*,1j}{\bm{X}})^{\top}{\bm{X}})\Big{[}\sum_{u_{1},v_{1}=1}^{d}\tilde{M}_{n,j,e_{u_{1}v_{1}}}{\bm{X}}^{(u_{1})}{\bm{X}}^{(v_{1})}$
$\hskip 113.81102pt+\sum_{u_{1},v_{1}=1}^{d}\sum_{u_{2},v_{2}=1}^{d}\tilde{M}_{n,j,e_{u_{1}v_{1}}}{\bm{X}}^{(u_{1})}{\bm{X}}^{(v_{1})}{\bm{X}}^{(u_{2})}{\bm{X}}^{(v_{2})}\Big{]}f_{G_{n}}({\bm{X}})$
$-\sum_{j=1}^{L}\tilde{N}_{n,j}\exp((W_{*,2}W_{*,1j}{\bm{X}})^{\top}{\bm{X}})f_{G_{n}}({\bm{X}})$
$+\sum_{j=1}^{L}\tilde{N}_{n,j}\exp((W_{*,2}W_{*,1j}{\bm{X}})^{\top}{\bm{X}})W_{*,2}W_{*,1j}{\bm{X}}$
$+\hat{R}_{n,1}({\bm{X}})+\hat{R}_{n,2}({\bm{X}})-\tilde{R}_{n,3}({\bm{X}})-\tilde{R}_{n,4}({\bm{X}})$ (34)
where $\tilde{N}_{n,j}:=\sum_{i\in\mathcal{V}_{j}}\exp(b_{n,i})-\exp(b_{*,j})$ for any $j\in[L]$.
@@
An important insight from equation (34) is that the ratio $Q_{n}({\bm{X}})/\mathcal{D}_{1n}$ can be expressed as a linear combination of the following independent functions:
@@
We proceed to demonstrate that not all of the coefficients of these linear independent functions go to 0 as $n$ goes to infinity. We prove that claim by contradiction. We assume that all these coefficients go to 0 as $n$ goes to $\infty$. From the representation of the ratio $Q_{n}({\bm{X}})/\mathcal{D}_{2n}$ in terms of these linear independent functions, we obtain that these ratios $\tilde{M}_{n,j,\alpha}/\mathcal{D}_{2n}$, $\tilde{M}_{n,j,\alpha,\beta}/\mathcal{D}_{2n}$, and $\tilde{N}_{n,j}/\mathcal{D}_{2n}$ approach 0 as $n\to\infty$, for all $\alpha,\beta\in\mathbb{N}^{d\times d}$ such that $1\leq|\alpha|+|\beta|\leq 2$.
By first considering the vanishing of the ratio $\tilde{N}_{n,j}/\mathcal{D}_{2n}\to 0$, we find that
$\frac{|\sum_{i\in\mathcal{V}_{j}}\exp(b_{n,i})-\exp(b_{*,j})|}{\mathcal{D}_{2n}}=\frac{|\tilde{N}_{n,j}|}{\mathcal{D}_{2n}}\to 0,$
for any $1\leq j\leq L$. By varying the index $j$ from 1 to $L$ in these limits and summing these limits, we find that
@@
Our strategy now is to consider the limits corresponding to Vornoi cells with exactly one element and more than one element separately. In particular, for Voronoi cells with exactly one element, namely, for indices $j\in[L]$ such that their corresponding Voronoi cells have one element, i.e., $|\mathcal{V}_{j}|=1$, as the ratio $\tilde{M}_{n,j,e_{uv}}/\mathcal{D}_{2n}\to 0$, we have
$\frac{\sum_{i\in\mathcal{V}_{j}}\exp(b_{n,i})\|\Delta W_{n,2}W_{n,1ij}\|_{1}}{\mathcal{D}_{2n}}=\frac{\sum_{u,v=1}^{d}|\tilde{M}_{n,j,e_{uv}}|}{\mathcal{D}_{2n}}\to 0.$
@@
We now move to the Voronoi cells with more than one element, namely, indices $j\in[L]$ such that $|\mathcal{V}_{j}|>1$. As the ratio $\tilde{M}_{n,j,2e_{uv}}/\mathcal{D}_{2n}\to 0$, we obtain that
$\frac{\sum_{i\in\mathcal{V}_{j}}\exp(b_{n,i})\|\Delta W_{n,2}W_{n,1ij}\|^{2}}{\mathcal{D}_{2n}}=\frac{\sum_{u,v=1}^{d}\tilde{M}_{n,j,2e_{uv}}}{\mathcal{D}_{2n}}\to 0.$ (37)
Combining all the results in equations (35), (36), and (37), we obtain that
$\frac{\mathcal{D}_{2n}}{\mathcal{D}_{2n}}=1\to 0,\mathrm{as}\ n\to\infty,$
which is a contradiction. As a consequence, at least one of the coefficients of the linear independent functions in the expression of the ratio $Q_{n}({\bm{X}})/\mathcal{D}_{2n}$ does not go to 0 as $n$ approaches infinity.
@@
The idea of this step is to divide all of the coefficients of the linear independent terms in the expression of the ratio $Q_{n}({\bm{X}})/\mathcal{D}_{2n}$, namely, the terms $\tilde{M}_{n,j,\alpha}/\mathcal{D}_{2n}$, $\tilde{M}_{n,j,\alpha,\beta}/\mathcal{D}_{2n}$, and $\tilde{N}_{n,j}/\mathcal{D}_{2n}$ for all $\alpha,\beta\in\mathbb{N}^{d\times d}$ such that $1\leq|\alpha|+|\beta|\leq 2$, by the maximum of their absolute values. In particular, we first denote denote $m_{n}$ as the maximum of the absolute values of those coefficients. As not all of these coefficients go to 0, it demonstrates that $1/m_{n}$ does not go to infinity as $n\to\infty$.
From the hypothesis, we have $\|f_{G_{n}}-f_{\widetilde{G}_{*}}\|_{L_{2}(\mu)}/\mathcal{D}_{2n}\to 0$ as $n\to\infty$. Since $1/m_{n}\not\to\infty$, it follows that $\|f_{G_{n}}-f_{\widetilde{G}_{*}}\|_{L_{2}(\mu)}/(m_{n}\mathcal{D}_{2n})\to 0.$ An application of the Fatou’s lemma leads to
$0=\lim_{n\to\infty}\dfrac{\|f_{G_{n}}-f_{\widetilde{G}_{*}}\|_{L_{2}(\mu)}}{m_{n}\mathcal{D}_{2n}}\geq\int\liminf_{n\to\infty}\dfrac{\left|f_{G_{n}}({\bm{X}})-f_{\widetilde{G}_{*}}({\bm{X}})\right|}{m_{n}\mathcal{D}_{2n}}d\mu({\bm{X}})\geq 0$
Combining all of the above results, we obtain that $\liminf_{n\to\infty}\dfrac{\left|f_{G_{n}}({\bm{X}})-f_{\widetilde{G}_{*}}({\bm{X}})\right|}{m_{n}\mathcal{D}_{2n}}=0$ for almost surely ${\bm{X}}$. To ease the presentation, we denote the following limits:
$\dfrac{\tilde{M}_{n,j,\alpha}}{m_{n}\mathcal{D}_{2n}}\to\lambda_{j,\alpha},\quad\dfrac{\tilde{M}_{n,j,\alpha,\beta}}{m_{n}\mathcal{D}_{2n}}\to\xi_{j,\alpha,\beta},\quad\dfrac{\tilde{N}_{n,j}}{m_{n}\mathcal{D}_{2n}}\to\tau_{j},$
for any $1\leq j\leq L$ and $\alpha,\beta\in\mathbb{N}^{d\times d}$ such that $1\leq|\alpha|+|\beta|\leq 2$. By the definition of $m_{n}$, at least one coefficient in the set $\{\lambda_{j,\alpha},\xi_{j,\alpha,\beta},\tau_{j}:j\in[L],\alpha,\beta\in\mathbb{N}^{d\times d}:1\leq|\alpha|+|\beta|\leq 2\}$ must be nonzero. Then, the equation $\liminf_{n\to\infty}\dfrac{\left|f_{G_{n}}({\bm{X}})-f_{\widetilde{G}_{*}}({\bm{X}})\right|}{m_{n}\mathcal{D}_{2n}}=0$, or equivalently, $\liminf_{n\to\infty}\dfrac{\left|Q_{n}({\bm{X}})\right|}{m_{n}\mathcal{D}_{2n}}=0$ leads to
@@
$\lim_{\varepsilon\to 0}\inf_{G\in\mathcal{G}_{L^{\prime}}(\Theta):\mathcal{D}_{2}(G,\widetilde{G}_{*})\leq\varepsilon}\|f_{G}-f_{\widetilde{G}_{*}}\|_{L_{2}(\mu)}/\mathcal{D}_{2}(G,\widetilde{G}_{*})>0,$
which suggests that the conclusion of the local part of the inequality (32) is proved.
B.0.2 Global part.
@@
From the result of the local part of the inequality (32), we can find a positive constant $\varepsilon^{\prime}$ such that the following inequality holds:
$\inf_{G\in\mathcal{G}_{L^{\prime}}(\Theta):\mathcal{D}_{2}(G,\widetilde{G}_{*})\leq\varepsilon^{\prime}}\|f_{G}-f_{\widetilde{G}_{*}}\|_{L_{2}(\mu)}/\mathcal{D}_{2}(G,\widetilde{G}_{*})>0.$
To obtain the conclusion of the theorem, we only need to demonstrate that
$\inf_{G\in\mathcal{G}_{L^{\prime}}(\Theta):\mathcal{D}_{2}(G,\widetilde{G}_{*})>\varepsilon^{\prime}}\|f_{G}-f_{\widetilde{G}_{*}}\|_{L_{2}(\mu)}/\mathcal{D}_{2}(G,\widetilde{G}_{*})>0.$
We prove the claim by contradiction. Indeed, by assuming the claim does not hold, it implies that there exists a sequence of $G^{\prime}_{n}:=\sum_{j^{\prime}=1}^{L^{\prime}}\exp(b_{n,j^{\prime}})\delta_{W_{n,1j^{\prime}}W_{n,2}}$ in the set $\mathcal{G}_{L^{\prime}}(\Theta)$ such that